OpenStudy (anonymous):

What is the standard form equation of the line passing through (3, –2) with a slope of 3/5?

OpenStudy (phoenixfire):

\[y=mx+c\]You're given the slope m=3/5, and a point (x, y)=(3, -2). So substitute the point and slope into that equation and solve for C to find your intercept.

OpenStudy (anonymous):

19/5

OpenStudy (anonymous):

so that would be....

OpenStudy (anonymous):

5x – 3y = 19

OpenStudy (phoenixfire):

\[c=-{19 \over 5}\] So your standard form is\[y={3 \over 5}x-{19 \over 5}\]

OpenStudy (anonymous):

then multiply everything by 5 and your good. thats right.

OpenStudy (anonymous):

thank you